OpenStudy (anonymous):

Find the roots 2x*2 + 4x - 5 = 0

OpenStudy (diyadiya):

Do you know the quadratic formula?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i just need to know the answers because i need to hury up and get my homework turned in

OpenStudy (anonymous):

wolfram buddy

OpenStudy (anonymous):

http://www.wolframalpha.com/

OpenStudy (diyadiya):

Sorry! Please read the Code of Conduct Its against the rules of Openstudy to give 'Just the answers' http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

type in the problems and it will give you the answers

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

it didnt work for that question

OpenStudy (blues):

Guys, OpenStudy is a tutoring service and not an answering site. The aim is to teach, not do other peoples' homework but to help them do it themselves. And big part of that is being nice to others.

OpenStudy (anonymous):

ok blues can you help me

OpenStudy (blues):

Though I am not a math person I will certainly try. The roots of the equation are the values of x, at which the entire equation equals zero. The easiest way to find them is to factor the equation. That means put it in (x + a number)*(x + another number) = 0 form.

OpenStudy (blues):

Or, because I don't think this one factors easily, use the quadratic formula...

OpenStudy (anonymous):

;p

OpenStudy (diyadiya):

Yes ,and you can use the quadratic equation to find the roots \[\Large x= \frac{-b \pm \sqrt{b^2-4ac}}{2a}\] Here a =2 ,b=4 ,c=-5

OpenStudy (anonymous):

so what would the answer be to that

OpenStudy (anonymous):

x=5/8

OpenStudy (diyadiya):

Substitute the values of a,b & c to the formula you'll get it!